### v1

# Dicke-Ising quantum battery of an ion chain driven by a mechanical oscillator

Abstract
A scheme for implementing quantum batteries in a realizable and controllable platform based on a trapped ion chain driven by a mechanical oscillator is proposed. The effects of the hopping interaction between the two-level ions and the coupling interaction between the ions and the external mechanical oscillator on the charging process of the battery are investigated. The importance of the counter-rotating wave terms in the system’s Hamiltonian, which are often ignored, is analyzed, and it is found that the charging energy and the ergotropy of the battery are dramatically affected by the counter-rotating wave terms. The quantum phase transition of the two-level system is restrained by the counter-rotating wave terms due to the destruction of the quantum coherence. Lastly, the power-law dependence of the charging process on the distance between the ions is discussed. Our theoretical analysis provides a solid foundation for the development of a practical quantum battery.
I I. Introduction
The battery is an important device that can store, deliver and convert energy. Such device has produced a huge impact on modern civilization and attracted great enthusiasm for researches since its invention. With the development of the science and technology, more and more researchers have been trying to apply the quantum theories and techniques to the battery for improving its performance. In the process of the development, the concept of the quantum battery was put forward F.Campaioli2023 . In 2013, R. Alicki and M. Fannes built the fundamental theory of the quantum battery’s charging and investigated the relationship between the passive states, the ergotropy and the Gibbs canonical density matrix RAMF . In the same year, a method at the price of time to extract all possible work without creating any entanglement was provided, and the relation between the entanglement and the ergotropy was discussed KVHM . In 2015, the protocol for charging the single qubit was demonstrated and then extended to the system containing $N$ qubits FCBS . F. Campaioli et.al pointed out that the charging power would be enhanced if $N$ batteries are charged collectively FCFA . In 2018, it has been found that the anisotropic interactions for the spin chain can provide a boost to the charging power, but the analysis was done under the condition of the weak interaction TPLJ . G. M. Andolina et.al studied the reason of the fast charging of quantum batteries, including the effects of many-body interaction and quantum-mechanical nature of the model itself, the result indicated that it depends on the specific models and parameters GMAM . Optimal control methods are used for preparing quantum states and recently they are exploited for fast charging quantum batteries F.Mazzoncini2023 ; RRRodriguez2024 . In addition, the realistic quantum system is affected by environment inevitably, and an increasing number of research works are devoting to the study of quantum batteries under the framework of open systems JMMF ; FHKF ; SYBJ ; FZFQ ; RRRB ; Downing2023 . The explorations above and other related works have laid foundation for the further investigation of the quantum batteries.
Inspired by the recent progress, a question arises: is there a realistic platform, which is easy to be prepared and manipulated, to develop the quantum batteries? At present, several schemes of quantum batteries based on superconductors CKHu2022 , quantum dots Wenniger2022 , organic microcavities JQQuach2022 and nuclear spin JJoshi2022 have been proposed, but they are still far from practical applications. The ion trap, which has been thought as an excellent platform, has demonstrated many important applications in the quantum fields, including quantum phase transition XLDD ; DPJI , quantum simulation RBCF ; Duan2024 and quantum computation KKim2019 ; C.Noel2022 . In particular, the study done by Duan and coworkers has realized a quartic trapping potential and allowed one to trap and control 512 ions experimentally Duan2024 . How to realize the quantum batteries in the platform of ion trap? The technique of the quantum optics allows us to regulate and control the trapped ions accurately XLDD , which is the basis of realizing the quantum batteries of trapped ion chains.
In the present paper, a practical proposal to implement the quantum battery based on a trapped ion chain is shown. The ions trapped in a Paul trap are coupled with a mechanical harmonic oscillator, as shown in Fig. 1. The pump field is used to drive the oscillator. The coupling between the mechanical mode and the trapped two-level ions due to the gradient magnetic field is tunable precisely PBLY . The magnetic field is produced by a magnetic film deposited on a cantilever with nanoscale thickness. The system can transfer the energy of phonons to the quantum battery by the interactions of the spin and mechanical degree of freedom. The $N$ two-level ions are assumed to couple the external mechanical oscillator with same strength. We investigate the charging process of the trapped ion quantum battery affected by the hopping interaction between the ions and the coupling interaction between the ions and the mechanical oscillator, discuss the effect of the counter-rotating wave terms of the system’s Hamiltonian on the process, and obtain the maxima of the charging energy and ergotropy under certain situations. It is hoped that our research can provide a reference for the development of the realistic quantum batteries.
The rest parts of the paper are structured as follows. In Sec. II, we introduce the Hamiltonian of our charging system, define some important physical quantities to characterize the charging process of the system. Sec. III gives the numerical results and explains the physics behind the phenomena. A brief conclusion is given in the last section.
Figure 1: (Color online) Schematic diagram of a quantum battery based on a trapped ion chain coupled with a mechanical harmonic oscillator (the cantilever that is fixed at one end and suspended at the other). The pump field is used to drive the oscillator and the coupling parameter $\lambda$ between the ion chain and the oscillator is tunable precisely. The green region represents the trapping field for the ion train.
II II. Model and characterizing method
Fig. 1 is the sketch of our charging system. An ion chain with $N$ two-level ions, in which the level splitting of each ion is $\omega_{a}$, is coupled by a mechanical harmonic oscillator with the angular frequency $\omega_{c}$. The coupling strength between the chain and the oscillator is presented by a controllable parameter $\lambda$. The hopping interaction between the different ions is proportional to $J/\left|z_{m}-z_{n}\right|^{p}$, in which $z_{m}$ and $z_{n}$ are the scaled equilibrium positions for the $m$th and $n$th ions, respectively. The value of the tunable exponent $p$ is chosen at the range $[0,3]$ XLDD ; PJBP . The Hamiltonian of the system is
$\hat{H}$ $=$ $\hat{H}_{c}+\hat{H}_{a}+\hat{H}_{ac},$
$\hat{H}_{a}$ $=$ $\omega_{a}\sum_{n=1}^{N}\hat{\sigma}_{n}^{+}\hat{\sigma}_{n}^{-}+J\sum_{n=1}^{N}\sum_{m>n}^{N}\frac{\hat{\sigma}_{n}^{x}\hat{\sigma}_{m}^{x}}{\left|z_{m}-z_{n}\right|^{p}},$
$\hat{H}_{c}$ $=$ $\omega_{c}\hat{c}^{+}\hat{c},~{}~{}\hat{H}_{ac}=\lambda\sum_{n=1}^{N}(\hat{c}+\hat{c}^{+})\hat{\sigma}_{n}^{x}.$ (1)
Here, $\hat{\sigma}_{n}^{+}\hat{\sigma}_{n}^{-}=(I_{n}+\hat{\sigma}_{n}^{z})/2$ with $\hat{\sigma}_{n}^{\pm}=(\hat{\sigma}_{n}^{x}\pm i\hat{\sigma}_{n}^{y})/2$. $\hat{\sigma}_{n}^{\alpha}$ and $I_{n}$ are the Pauli and identity matrices for the $n$th ion, where $n=1,2,...,N$ and $\alpha=x,y,z$. $\hat{H}_{c}$ is the Hamiltonian of the mechanical harmonic oscillator, $\hat{c}$ ($\hat{c}^{+}$) is the corresponding annihilation (creation) operator. $\hat{H}_{a}$ includes their own energy of the two-level ions and the hopping energy between the different ions. $\hat{H}_{ac}$ represents the coupling interaction between the trapped ion chain and the oscillator. The Hamiltonian $\hat{H}$ in Eq. (II) can be regarded as the Dicke-Ising model. We have set $\hbar\equiv 1$ in the above Hamiltonian. For convenience but without losing the physics, the parameter $\omega_{c}=\omega_{a}=1.0$ is chosen.
We consider the charging process of the system which consists of the mechanical harmonic oscillator and five $(N=5)$ two-level ions under the view of the closed system. Our aim is to construct a charging scheme of the quantum battery which can collect low-frequency vibration energy. The influence of the hopping interaction between the ions and the coupling interaction between the ions and the oscillator on the charging process is investigated, with a particular focus on the significance of the counter-rotating wave terms associated with these two interactions. The initial state of the two-level ions is the ground state $\left|g\right\rangle_{a}$ obtained by numerically diagonalizing $\hat{H}_{a}$. The mechanical oscillator is assumed to be prepared in the superposition of Fock states $\left|\Phi\right\rangle_{c}=\sqrt{0.6}\left|10\right\rangle+\sqrt{0.4}\left|15\right\rangle$. There have been a lot of theoretical studies that have suggested synthesizing superpositions of Fock states in different systems VBergholm2019 ; PDNation2013 ; QZYY ; MKounalakis2019 ; H.Tan2014 ; H.Xie2019 ; JJing2024 . Experimentally, the multi-phonon Fock states and their superposition states have only been obtained inside a high-overtone bulk acoustic wave resonator as far as we know Y.Chu2018 . In the following context, we assume that the initial wavefunction of such coupling system is $\left|\Psi(0)\right\rangle=\left|\Phi\right\rangle_{c}\otimes$ $\left|g\right\rangle_{a}$, and the scaled equilibrium positions $z_{n}$ of the five trapped ions are $\left[-1.7429,-0.8221,0,0.8221,1.7429\right]$ in the Paul trap DFVJ . The dynamical evolution for the system is based on
$\left|\Psi(t)\right\rangle=e^{-i\hat{H}t}\left|\Psi(0)\right\rangle,$ (2)
and $\hat{\varrho}(t)=\left|\Psi(t)\right\rangle\left\langle\Psi(t)\right|$. The reduced density operator is $\hat{\varrho}_{a}(t)=Tr_{c}[\hat{\varrho}(t)]$ and the subscript $c$ indicates a partial trace. The total energy of the ion chain can be calculated by $E(t)=Tr[\hat{H}_{a}\hat{\varrho}_{a}(t)]$. The energy obtained from the mechanical oscillator is $E_{c}(t)=E(t)-E(0)$, which represents the actual charging energy of the battery. Another key quantity is ergotropy defined as $E_{e}(t)=E(t)-\sum_{k}r_{k}(t)e_{k}$ Allahverdyan2004 , and the eigenvalues $r_{k}(t)$ and $e_{k}$ are determined by
[…]
Here, $r_{k}(t)$ are arranged in descending order but $e_{k}$ in ascending order. The ergotropy is used to describe the maximal extractable work from the ion chain. The dynamics of the model (II) can not be obtained analytically. The Hilbert space of the quantum harmonic oscillator must be cut off for numerical calculations. The dimension of the space is set to $101$ after the space has been truncated, and the eigenstate for the highest energy level is $\left|100\right\rangle$. The dimension number of the Hilbert space exceeds four times of the phonon’s number for the harmonic oscillator DFMC .
Several other quantities are introduced to characterize the charging process of the system. $\sigma_{n}$ represents the expectation value of $\hat{\sigma}_{n}^{+}\hat{\sigma}_{n}^{-}$, which is defined as $\left\langle\hat{\sigma}_{n}^{+}\hat{\sigma}_{n}^{-}\right\rangle$, and $\sigma_{m,n}=\sigma_{m}-\sigma_{n}.$ In order to show the quantum phase transition, we need to introduce $M_{z}=\left\langle g\right|\hat{S}_{z}\left|g\right\rangle_{a}/N$ and $O_{z}=\left\langle g\right|\hat{S}_{z}^{2}\left|g\right\rangle_{a}/N^{2}$, where $S_{z}=\sum_{n=1}^{N}\sigma_{n}^{z}$ FQDQ . The entropy is significant in both classical and quantum physics, and the von Neumann entropy is used to quantify entanglement of the pure states CHBD . According to the definition, the von Neumann entropy is expressed as $S(t)=-\sum_{k}r_{k}(t)\log_{2}r_{k}(t)$. It generally requires the dimensions of the two subsystems are same, but our model does not meet this condition. The calculated values of the von Neumann entropy for the two subsystems are equal. So we can use the von Neumann entropy to describe the disorder of the subsystem and the degree of entanglement for the whole system.
III III. Results and their physical interpretations
Figure 2: (Color online) The charging process of the quantum battery in our model. (a) The evolution of the charging energy $E_{c}(t)$ and the ergotropy $E_{e}(t)$; (b) The evolution of the entropy $S(t)$; (c) The dynamics of the expectation value $\left\langle\hat{\sigma}_{3}^{+}\hat{\sigma}_{3}^{-}\right\rangle$ of the third ion; (d) The difference of the expectation value between the different ions. The parameters have been set as $\lambda=0.25$, $J=0.2$ and $p=3$.
To guarantee that the quantum battery on our platform receives sufficient energy, we have assumed that the number of the phonons is 12, which is greater than two times of the number of the ions. Actually, the charging process of the quantum batteries is dynamical one, which can be investigated based on Eqs. (II) and (2). The quantities for characterizing the process are determined by numerically solving the above equations. Fig. 2 displays the dynamical evolution of the charging energy $E_{c}(t)$, the ergotropy $E_{e}(t)$ and the entropy $S(t)$. In Fig. 2(a), the evolution of $E_{c}(t)$ and $E_{e}(t)$ gives an irregular oscillation with a small amplitude around an average value after experiencing a relaxation process. The difference between the local maxima and minima (except for $t=0$) for $E_{c,e}(t)$ is relatively small. Such results also hold for the time range $\omega_{c}t\in[0,100]$. They are very different from the results in Ref. FQDQ since there exists beating structure in the evolution of $E_{c,e}(t)$ therein. In addition, the difference between the local maxima and minima (except for $t=0$) of $E_{c,e}(t)$ are larger than ours. Both of the counter-rotating wave terms and the non-nearest-neighbour hopping have been considered in our model. For $J=0.2$, the effect of the nearest-neighbor hopping is relatively small, and the effect of the non-nearest-neighbour hopping can be ignored. The main reason for the above results is due to the role played by the counter-rotating wave term $\hat{H}_{ac,cw}\equiv$ $\lambda\sum_{n=1}^{N}(\hat{c}^{+}\hat{\sigma}_{n}^{+}+\hat{c}\hat{\sigma}_{n}^{-})$ in the Hamiltonian $\hat{H}_{ac}$ for the coupling interaction between the ion chain and the oscillator.
Maybe, some people have a doubt that whether the counter-rotating wave term can play an important role in our system because of $\lambda/\omega_{c}=0.25$ in Fig. 2? In fact, the counter-rotating wave term must be taken into account for $\lambda/\omega_{c}\geq 0.1$ since the system has entered the so-called ultrastrong coupling regime XTZH . Previous studies have demonstrated that the counter-rotating wave term can induce the collapse and revival of the quantum dynamics JCGR ; JFHC . In 2008, H. Zheng et.al highlighted that the Zeno time is longer by $2$ orders of magnitude than that calculated without considering the counter-rotating wave term HZSY . Here, the counter-rotating wave term destroys the quantum coherence and leads to the disappear of the beating structure, and finally restrains the energy exchange between the two-level ions and the mechanical oscillator. In Fig. 2(b), the entropy $S(t)$ rapidly increases from the initial time and then approximately approaches to a stable value. This means the entanglement between the ion chain and the external driving mechanical oscillator is stabilized with the increase of charging time. As an important quantum resource, the change of the entanglement does not show a positive correlation with $E_{c}(t)$ and $E_{e}(t)$ during the charging process. The dynamics of $\sigma_{3}$, which corresponds to the average energy of the most intermediate ion in the chain, is demonstrated in Fig. 2(c). $\sigma_{m,n}$ indicates that there is the energy difference of the $m$th and $n$th ions although they couple the mechanical oscillator with the same strength, as shown in Fig.2 (d). It indicates that the contribution to the charging process of every ion is different, and the third ion is not superior or inferior although it has a special position in the ion chain.
Figure 3: (Color online) The comparison of the charging process for the charging energy $E_{c}(t)$ in (a) and the ergotropy $E_{e}(t)$ in (b) under different coupling strength $\lambda$. The other parameters are set as $J=0.4$ and $p=3$.
Fig. 3(a) and (b) show the evolution of $E_{c}(t)$ and $E_{e}(t)$ under different coupling strength $\lambda$, respectively. Similar to the result for $\lambda=0.25$ and $J=0.2$ given in Fig. 2(a), there exist some irregular oscillations after relaxation at the initial stage. The oscillation amplitude of $E_{c,e}(t)$ is larger than the other two cases corresponding to $\lambda=0.5$ and $\lambda=1.0$ in Fig. 3(a) since the role played by the counter-rotating wave term $\hat{H}_{ac,cw}$ is relatively small for $\lambda=0.1$. The counter-rotating wave term $\hat{H}_{ac,cw}$ becomes more and more important with increasing $\lambda$. The local maxima of $E_{c}(t)$ and $E_{e}(t)$ have been suppressed heavily for larger $\lambda$. The speed of energy exchange in the process is increased by increasing the coupling strength $\lambda$ and then reducing the oscillation period, which leads to the shorter time to reach the first local maximum. The stable value of $S(t)$ after a long period of time is improved by increasing $\lambda$. The quantum entanglement between the ion chain and the mechanical oscillator is enhanced. But it has no promoting effects for the maximization of $E_{c}(t)$ and $E_{e}(t)$ DLYang2023 . From the view of the subsystem, the disorder of the two-level system can be described by the entropy $S(t)$. According to the classical thermodynamics, the disorder energy is more difficult to convert into the available work, so that the ergotropy $E_{e}(t)$ decreases with increasing $\lambda$ shown in Fig. 3(b) can be explained. If the charging time is defined by the time at which $E_{e}(t)$ firstly reaches its local maximum, the charging becomes quicker for larger $\lambda$. So increasing $\lambda$ can improve the charging speed but suppress the ergotropy. A reasonable strength of the coupling interaction between the ions and the oscillator should be taken for the actual battery platforms.
Figure 4: (Color online) The comparison of the charging process for the charging energy $E_{c}(t)$ in (a) and the ergotropy $E_{e}(t)$ in (b) under different strength of the hopping interaction $J$. (c) and (d) present the corresponding results without considering the counter-rotating wave term between the two-level ions. The other parameters are set as $\lambda=0.5$ and $p=3$.
Fig. 4(a) and (b) show the time evolution of $E_{c}(t)$ and $E_{e}(t)$ under different $J$, respectively. According to the result in Fig. 4(a) , the final value of $E_{c}(t)$ exceeds about $5\omega_{a}$ (the highest energy of five two-level ions without considering the hopping energy) for larger $J$ after the initial relaxation. As increasing the hopping interaction, the ground energy becomes more lower and the energy spectrum is expanded dramatically [as shown in Fig. 6(a)], which is beneficial for charging and storing more energy in the quantum battery. But for small value of $J=0.1$, the final value of $E_{c}(t)$ is smaller that $3\omega_{a}$. The reason is that only a subset of the ions are excited in this situation, and the contribution from the hopping interaction is insufficient for small $J$. In Fig. 4(b), the final value of $E_{e}(t)$ is reduced when $J$ is greater than a critical value for the specific range. That is to say, the final value is not monotonic with increasing $J$. The similar result can also be obtained for $E_{c}(t)$ when $J=1.0$ and $J=2.0$ from Fig. 4(a). This can be interpreted by the enhancement of the counter-rotating wave term $\hat{H}_{J,cw}\equiv J\sum_{n=1}^{N}\sum_{m>n}^{N}(\hat{\sigma}_{n}^{+}\hat{\sigma}_{m}^{+}+\hat{\sigma}_{n}^{-}\hat{\sigma}_{m}^{-})/\left|z_{m}-z_{n}\right|^{3}$ between the ions. Fig. 4(c) and (d) show the results under the same conditions as in Fig. 4(a) and (b), but without considering the role of the counter-rotating wave term $\hat{H}_{J,cw}$. The final values of $E_{c}(t)$ and $E_{e}(t)$ increase with increasing $J$ in Fig. 4(c) and (d). As a higher-order nonlinear effect, the counter-rotating wave term $\hat{H}_{J,cw}$ between the ions plays an important role in the charging process of the quantum battery.
The maxima of $E_{c}(t)$ and $E_{e}(t)$ at the time range $\omega_{c}t\in[0,30]$ are not monotonic functions of $J$ and $\lambda$, as shown in Fig. 5. The charging energy is always larger than the ergotropy of the battery. The dependence of the maxima of $E_{c}(t)$ and $E_{e}(t)$ on $\lambda$ is given in Fig. 5(a). The peak is at $\lambda\approx 0.2$, which is the balance point of the competition between the rotating wave term and the counter-rotating wave one in the Hamiltonian $\hat{H}_{ac}$. The counter-rotating wave term $\hat{H}_{ac,cw}$ plays less important role in the charging process for $\lambda<0.2$. That is why in most of the references the rotating wave approximation has been used for studying the quantum dynamics when the value of $\lambda$ is small. Generally, the maxima of $E_{c}(t)$ and $E_{e}(t)$ decrease with increasing $\lambda$ for $\lambda>0.2$. The counter-rotating wave term $\hat{H}_{ac,cw}$ between the ion chain and the mechanical oscillator becomes dominant for larger $\lambda$, which is not conducive for the battery’s charging. The dependence of the maxima of $E_{c}(t)$ and $E_{e}(t)$ on $J$ is given in Fig. 5(b). The change of maxima of $E_{c}(t)$ and $E_{e}(t)$ with $J$ have several extremum points in Fig. 5(b). Increasing the hopping strength $J$ reduces the low energy levels and expands the whole energy spectrum, but improves the role of the counter-rotating wave term $\hat{H}_{J,cw}$ of the hopping interaction. The former is helpful for the storage of the energy in the ion chain, but the latter breaks the quantum coherence and suppresses the charging. The extremum points come from the competition between both of them in the charging process.
Figure 5: (Color online) The change of the maximum of the charging energy $E_{c}(t)$ and the ergotropy $E_{e}(t)$ with the coupling strength $\lambda$ in (a) for $J=0.3$ and the strength of the hopping interaction $J$ in (b) for $\lambda=0.4$. The time range of the calculation is $\omega_{c}t\in[0,30]$ and the exponent $p=3$ is used.
Figure 6: (Color online) The change of the eigenvalues $e_{k}$ for $\hat{H}_{a}$ with $J$ in (a, c) and $M_{z}$ and $O_{z}$ in (b, d). The counter-rotating wave term $H_{J,cw}$ is considered in (a, b), but not in (c, d). The exponent $p=3$ is used for both of the cases.
A quantum phase transition is supposed to occur in the thermodynamic limit of an infinite number of spins. In Ref. AFHS , the authors have pointed out that there are no quantum phase transition for the system with finite two-level ions by only considering the nearest-neighbor hopping interaction between the ions. Fig. 6(a) and (b) show the changes of energy spectrum of the ion chain, $O_{z}$ and $M_{z}$ with $J$ for considering the counter-rotating wave term $\hat{H}_{J,cw}$. It is shown that there is no quantum phase transition although the nearest- and non-nearest-neighbor hopping interactions between the five ions are all considered in our model. Fig. 6(c) and (d) show the change of the energy spectrum, $O_{z}$ and $M_{z}$ with $J$ without considering the counter-rotating wave term $\hat{H}_{J,cw}$. The result about the quantum phase transition is consistent with that given in Ref. FQDQ , which ignores the counter-rotating wave term in the Su-Schrieffer-Heeger quantum battery. That is to say, the counter-rotating wave term between the two-level ions suppresses the quantum coherence and leads to the disappearance of the quantum phase transition.
Figure 7: (Color online) The charging process of the quantum battery for $p=0$, $1$, $2$, and $2.5$ in the hopping interaction. The evolutions of the charging energy $E_{c}(t)$ in (a), the ergotropy $E_{e}(t)$ in (b), the entropy $S(t)$ in (c), and the difference of the expectation value between the first and third ions $\sigma_{3,1}$ in (d) are given. The values of the other parameters are same with that in Fig. 2.
Furthermore, the influence of the non-nearest-neighbor hopping interaction for $E_{c}(t)$ and $E_{e}(t)$ is considered. When $J$ is relatively small $(J\ll\omega_{a,c})$, the effect of the non-nearest-neighbor hopping interaction can be ignored. However, if $J$ has a relatively large value, the non-nearest-neighbor interaction exerts a discernible influence on the charging process. In particular, the influence also works through the power-law dependence of the distance $\left|z_{m}-z_{n}\right|$ for the exponent $p$. The hopping interaction has infinite range for $p=0$ and is short-ranged for $p=3$. The power-law dependence in the hopping interaction between the ions is demonstrated in Fig. 7. By choosing different values of the exponent $p$, the evolutions of the charging energy $E_{c}(t)$, the ergotropy $E_{e}(t)$, the entropy $S(t)$, and the difference of the expectation value between the first and third ions $\sigma_{3,1}$ are given. When $p=0$, the system has the highest symmetry since the hopping interaction doesn’t depend on the distance and the time mean of $E_{c}(t)$ and $E_{e}(t)$ is larger than others with $p\neq 0$, as shown in Fig. 7(a) and (b). The evolution of $E_{c}(t)$ and $E_{e}(t)$ doesn’t change very much with the exponent $p$ when $p\geq 2$, showing that the change of $p$ has a little effect on the charging ability of the battery. The degree of the entanglement characterized by the final entropy is increased with increasing $p$ until $p\geq 2$ from Fig. 7(c). The dynamics of $\sigma_{3,1}$ in Fig. 7(d) has similar changing trends for different $p$ except $p=0$. That is to say, the hopping interaction between the ions exhibits the short-ranged behaviours for $p\geq 2$ and the result presented in Fig. 2 for $p=3$ serves as a representative example.
IV IV. Conclusion
In the present work, we have offered a proposal to realize the quantum batteries in the system of trapped ions, and pointed out that the energy of low-frequency vibration can be converted, stored and utilized. The effects of the hopping interaction between the two-level ions and the coupling interaction between the ions and the external mechanical oscillator on the charging process of the battery are mainly investigated. The complicated roles played by the counter-rotating wave terms in the ultrastrong coupling regime were revealed and discussed by the analytical and numerical calculations. It is found that: (i) The counter-rotating wave term in the coupling interaction between the ion chain and the oscillator enhances the quantum entanglement between the chain and the oscillator, improves the charging speed of the batteries, but suppresses the extractable ergotropy; (ii) The counter-rotating wave term between the ions destroys the quantum coherence, leads to the disappearance of the quantum phase transition, and suppresses the charging of the quantum batteries. The advantage of our model is that it does not ignore the counter-rotating wave terms and more related factors can be studied. The corresponding results can be directly used to analyze the situation without considering the rotating-wave approximation DBRA . Our work provides a scheme and numerical results for the implementation of realizable quantum batteries and will simulate the experimental research in this direction.